OpenStudy (anonymous):

Only 4 genius..... Other don't read... If 2+9=7 & 7+0=9 then 7-8=0 so what 2+1=?

OpenStudy (agent47):

2=-2 7-8=0, 7=8, 0=8-7=1, so 1=0 2+1=-2+0=-2

OpenStudy (agent47):

unless you're using some crazy made up algebraic rules

OpenStudy (anonymous):

u ansr is totly rong dude

OpenStudy (agent47):

*KRAZZy made up arithmetic rules

OpenStudy (anonymous):

6

OpenStudy (agent47):

Ok, then please state which numbers we should treat as numbers and which ones we should treat as variables

OpenStudy (anonymous):

0

OpenStudy (anonymous):

TWO + ONE = 6 letters

OpenStudy (agent47):

Hey guys no guesswork, please post your reasoning

OpenStudy (agent47):

mathwhiz is correct

OpenStudy (anonymous):

TWO + NINE = 7 letters SEVEN + ZERO = 9 letters SEVEN - EIGHT = 0 letters

OpenStudy (agent47):

GENIUS!

OpenStudy (anonymous):

just solve 2+9=7 hw comes then next @Agent47

OpenStudy (agent47):

2+9=7 Solving for 2 gives: 2=7-9=-2

OpenStudy (anonymous):

yes math is ryt :)

OpenStudy (anonymous):

@mathwhiz99 is kinda right.. XD

OpenStudy (agent47):

So 2=-2 7-8=0, so -(7-8)=0, or 8-7=0, or 1=0 2+1=-2+0=-2 You can't say I'm wrong either ;D

OpenStudy (agent47):

well maybe you can lol, this whole thing might not be consistent with this equation: 7+0=9

OpenStudy (anonymous):

yeah bt u r going on rong way @Agent47

OpenStudy (agent47):

lol yes, I understand how he did it, but I'm just bored so I'm just showing alternative solutions.. that give alternative answers ;)

OpenStudy (anonymous):

:D